c) NHA (aq) H2O(l) NH3(aq) H30 (aq). d) HNO3(aq) H20 (l) NO3 (aq) H30 (aq). 16. What are the Bronsted-Lowry acids in the following chemical reaction? HBr(aq) CH3COOH (aq) CH3C(OH 2 (aq) Br (aq) a) CH3COOH, CH3C(OH)2 b) CH3COOH, Br c) HBr, CH3COOH d) HBr, CH3C(OH)2 17. What is the strongest acid among the following? a) H2SO3 b) 2Seo3 c) H2SO d) H2Seo4 18. What is the Kw of pure water at 50.0ec, ifthe pH is 6.630? a) There is not enough information to calculate the Kw. b) 50 x 10 14 1.00 x 10 14 d) 13 x 10-14 e) 2.34 x 10 19. Calculate the concentration of H3o in a solution that contains 5.5 x 10 s MoH at 25 c. ldentify the solution as acidic, basic, or neutral. 9.2 x 10 M, acidic by 10 M, basic 8 x 10 c) 1.8 x 10 M, acidic d) 5.5 x 10 10 M, neutral e) 9.2 x 10 M, basicPlease answer 18 and 19! and show work
Answer #1

18)

pH = 6.630

use:

pH = -log [H3O+]

6.630 = -log [H3O+]

[H3O+] = 2.344*10^-7 M

SINCE water is neutral,

[OH-] = [H3O+]

So,

[OH-] = 2.344*10^-7 M

Now use:

Kw = [H+] [OH-]

= (2.344*10^-7)*(2.344*10^-7)

= 5.50*10^-14

Answer: b

19)

use:

[H3O+] = (1.0*10^-14)/[OH-]

= (1.0*10^-14) / (5.5*10^-5)

= 1.8*10^-10 M

Since [OH-] > [H+], it is basic

Answer: b